To connect to the meeting, please do the following: Time: Wednesday at 9:00 AM EST.
- To join via a Web Browser, go to the page  https://bluejeans.com/984025598.
- To join via Polycom room system go to the IP Address: 188.8.131.52 (bjn.vc) and enter the meeting ID: 984025598.
- To join via phone, use one of the following numbers and the Conference ID: 984025598.
- Dial Either +1 408 740 7256 or +1 888 240 2560 for US or Canada
- 8/9/21 Start Hot Check-out
New items for discussion:
Short term schedule update:
Reports by Group: CIS/EGG: assed for HV commissioning (See ABIL5170)
- Gun HV Processing under way (UITFList 1663); up to 182kV so far
- HV/Kr processing in progress: Processed to 235 kV in Kr; evaluating in vacuum now.
- Gun rebuild schedule
- Corrector installation & cable reroute mostly complete, turned over to DC power
- Chris Gould making nomenclature changes / Gary Croke & Michele Joyce making controls changes / Eric Forman to update Quick Reference Drawing / Shaun to take care of song sheets
- Changes made in UED workspace for names and field maps; ready to merge soon. (Thanks Gary!)
- Solenoid PS being installed
Instrumentation and Controls:
- BCM project software integration under way between Musson/Seaton and Sue Witherspoon
- Chris Norris to hook up viewers next week.
- Brian Freeman available for adViewer work for video camera for ITVM604A; some focusing still required.
- 6GHz > 1.5GHz Down-converter installed for bunch length system
- BLM locations being set 7/21 following this meeting
- FSD inputs for magnet and BCM being prepared
ME / Installation:
- Pb stacked; RCG required modification due to permanent shielding req delay.
- New designs for FC3 and FC4 with a great many thanks to Shaun for the design work and John/Xi for Pb arrangement and table construction!
- Songsheet being updated by Shaun; be sure to let him know when changes are made!
- Status of permanent shielding purchase?
- COO/ESAD/RSAD prepared; committee response in progress.
- EES work happening in earnest
- List of Engineering Jobs remaining
Still to do:
- Add Mu metal over bellows on wastewater line.
- Mount bottle rack in Labyrinth (Staged); remove gas bottles to rack.
- Unistrut mounting for cables - Neal will do when available
- Engineering work on beam line starting
- Pull Heliax from Spectrum analyzer/Control room >> Buncher
- 4th terminal system in Control room under assembly/rebuild
- Items from To-do list...
Engineering Other business
Reports on Open Action Items:
- Formal run plan (McCaughan) Strawman